# Unsupervised full-field Bayesian inference of orthotropic hyperelasticity from a single biaxial test: a myocardial case study

### v1

@@ Abstract
Cardiac muscle tissue exhibits highly non-linear hyperelastic and orthotropic material behavior during passive deformation. Fully capturing this behavior in traditional homogenized tissue testing requires the excitation of multiple deformation modes, i.e. combined triaxial shear tests and biaxial stretch tests. Inherently, such multimodal experimental protocols necessitate multiple tissue samples and extensive sample manipulations. Intrinsic inter-sample variability and manipulation-induced tissue damage might have an adverse effect on the inversely identified tissue behavior. In this work, we aim to overcome this gap by focusing our attention to the use of heterogeneous deformation profiles in a parameter estimation problem. More specifically, we adapt EUCLID, an unsupervised method for the automated discovery of constitutive models, towards the purpose of parameter identification for highly nonlinear, orthotropic constitutive models using a Bayesian inference approach and three-dimensional continuum elements. We showcase its strength to quantitatively infer – with varying noise levels – the material model parameters of synthetic myocardial tissue slabs from a single heterogeneous biaxial stretch test. This method shows good agreement with the ground-truth simulations and with corresponding credibility intervals. Our work highlights the potential for characterizing highly nonlinear and orthotropic material models from a single biaxial stretch test with uncertainty quantification.
@@
material model inference , full-field data , uncertainty quantification , anisotropic hyperelasticity , experimental tissue testing , Bayesian inference
@@ 1 Introduction
Soft biological tissues exhibit complex mechanical responses due to their fibrous, anisotropic microstructure. Accurately characterizing this behavior is essential for constructing predictive constitutive models that support simulation-driven design, diagnostics, and treatment planning (Peirlinck et al., 2021). However, unlike engineered materials, soft tissues are subject to considerable biological variability, both within a patient, a patient cohort, and experimental study (Reeps et al., 2012; Roccabianca et al., 2014; Budday et al., 2017; Matouš et al., 2017). This intra- and inter-sample variability is even more pronounced in the myocardium, where strong spatial gradients in myofiber orientation and laminar architecture govern regional mechanical function (Rohmer et al., 2007; Alexander J. Wilson, 2022). In this study we investigate the impact that the underlying microstructural organization variability has on the mechanical behavior of a single sample.
@@
Traditional approaches to constitutive modeling of such tissues typically rely on multiple mechanical tests across different deformation modes including uniaxial tension or compression, biaxial tension, or shear testing (Dokos et al., 2002; Sommer et al., 2015; Kakaletsis et al., 2021). These protocols often require combining data from different specimens and, as such, ignore and smooth out this intrinsic intra- and inter-sample variability and lead to a smoothened out, average response (Staber and Guilleminot, 2017). Moreover, these protocols involve strong assumptions, such as spatially homogeneous stress and strain fields within the region of interest (Holzapfel and Ogden, 2009; Guan et al., 2018; Martonová et al., 2024). Such assumptions can be problematic for living tissues, where spatial heterogeneity is the norm, and the mechanical response may depend on subtle differences in fiber orientation, geometry, or mounting. Moreover, invasive sample manipulations, such as cutting, clamping, and gluing, can alter the intrinsic mechanical behavior that these tests seek to measure (Fehervary et al., 2018; Alloisio et al., 2024).
@@
These challenges have prompted the development of single-sample, single-experiment methodologies that aim to infer full constitutive behavior from a single loading protocol. This shift has been enabled by advances in full-field deformation measurements, such as digital image correlation and imaging-based 3D kinematics (Wang et al., 2021; Jailin et al., 2024; Meng et al., 2025). When applied to samples undergoing spatially heterogeneous deformation, such data enable the inference of multiple material parameters simultaneously, reducing the need for multiple loading modes. Indeed, unsupervised identification techniques, such as the virtual fields method (Grédiac et al., 2006; Pierron and Grédiac, 2012), finite element updating (Kavanagh and Clough, 1971; Peirlinck et al., 2018; Elouneg et al., 2021), variational systems identification (Wang et al., 2021), the equilibrium gap method (Claire et al., 2004; Peyraut and Genet, 2025), and EUCLID (Flaschel et al., 2021; Thakolkaran et al., 2022; Joshi et al., 2022; Thakolkaran et al., 2025), demonstrate that heterogeneous strain states, once seen as a nuisance, can be leveraged as rich sources of information, obviating the need for homogenizing assumptions. In this study, we explore the validity of a simple biaxial tensile testing protocol on a single sample towards the identification of the underlying orthotropic behavior of a myocardial tissue slab.
@@
Yet, a key limitation persists in most current workflows: the lack of uncertainty quantification. Material parameters are frequently reported as point estimates, with little accounting for the inherent measurement noise or biological variability. In the context of myocardial modeling, where even adjacent samples can differ markedly in stiffness and fiber orientation (Dokos et al., 2002; Nielles-Vallespin et al., 2017), such deterministic inference may be misleading (Hauseux et al., 2018). Recent work has begun to address this gap through Bayesian approaches to inverse modeling, which offer not only best-fit estimates but also posterior distributions reflecting uncertainty in data and model form (Ozturk et al., 2021; Joshi et al., 2022; Aggarwal et al., 2023; Peshave et al., 2024; Anton et al., 2024; Taç et al., 2024; Linka et al., 2025). Within the unsupervised setting, Bayesian-EUCLID (Joshi et al., 2022) forms an interesting contribution towards the inference of myocardial tissue behavior from full-field displacement fields and force measurements. Joshi et al. demonstrated the feasibility of using Gibbs sampling for full-field inference in two-dimensional problems, where the material response could be inferred from a linear system of equations for the unknown material parameters in a rich constitutive feature library. In contrast, the setting of this study involves two distinct and compounding challenges: first, the underlying constitutive models are highly nonlinear (Costa et al., 2001; Holzapfel and Ogden, 2009; Martonová et al., 2024), with strongly cross-correlated parameters (Guan et al., 2018; Laita et al., 2025) that give rise to complex posterior landscapes; and second, the transition to three-dimensional full-field data. While the latter provides richer information, it also substantially increases the computational cost of evaluating the likelihood function. These factors independently and jointly limit the practical utility of Markov Chain Monte Carlo (MCMC) methods, including Gibbs sampling, due to the high cost of repeated likelihood evaluations and the poor mixing behavior in high-dimensional, correlated spaces. These limitations motivate us to extend the Bayesian EUCLID framework with more scalable inference strategies, such as Stochastic Variational Inference (SVI), which approximates posterior distributions through optimization rather than sampling, enabling tractable inference in settings with both complex models and large data volumes.
@@
Towards these goals, this study presents an unsupervised, full-field, Bayesian framework for the inference of orthotropic hyperelastic material parameters from a single heterogeneous biaxial tensile test. Our approach builds on the Bayesian-EUCLID framework (Joshi et al., 2022), extended via SVI (Thorat et al., 2025) to capture parametric uncertainty in a highly nonlinear, and a-priori chosen, constitutive model form. The myocardium is modeled using the well-established Holzapfel–Ogden constitutive law (Holzapfel and Ogden, 2009; Avazmohammadi et al., 2019), and heterogeneity in the deformation field is introduced either through microstructural variation (i.e. intrinsic fiber-sheet-interlaminar architectural heterogeneity), geometric tissue sample modification, or both. We demonstrate that the spatial richness of the heterogeneous displacement field combined with reaction force data enables robust parameter inference with uncertainty quantification from a single test. We additionally validate the inferred material models using unseen invariant values and assess sensitivity to displacement noise.
This work highlights the potential of full-field, unsupervised, and uncertainty-aware approaches to redefine how we efficiently characterize complex living tissues, taking into account minimal tissue availability and minimizing tissue manipulation.
@@ 2 Unsupervised Stochastic Variational constitutive parameter inference
Conventional approaches for characterizing myocardial material behavior typically rely on spatially averaged measurements and simplifying assumptions of spatially homogeneous microstructural organization. However, this intrinsic homogenization obscures the true spatial heterogeneity present within a tissue sample. To address this, we infer the constitutive parameter vector in an unsupervised setting from experimentally accessible data: (i) readily available global reaction forces, (ii) full-field deformation, and (iii) the microstructural organization across the slab. In this work, we assume digital volume correlation (DVC) data derived from concomitant through-thickness imaging, such as ultrasound (Navy et al., 2025), Magnetic Resonance Imaging (Estrada et al., 2020), or X-Ray Imaging (Davis et al., 2024) to obtain three-dimensional deformation fields. Microstructural maps in (iii) can be obtained via SHG or micro-CT (Alberini et al., 2024; Kakaletsis et al., 2021; Maes et al., 2022). Among established inverse strategies such as Finite Element Model Updating (FEMU) (Meijer et al., 1999), the Virtual Fields Method (VFM) (Grédiac et al., 2006), the Equilibrium Gap Method (EGM) (Claire et al., 2004; Roux and Hild, 2020), and Efficient Unsupervised Unsupervised Constitutive Law Identification (EUCLID) (Flaschel et al., 2021), we build on the latter. EUCLID formulates the inverse problem via the weak form of momentum balance (see Section 2.1). Here, we extend it to a fully three-dimensional setting with nonlinear orthotropic materials, and recast the classic deterministic parameter optimization into an unsupervised Bayesian inference problem. Leveraging Stochastic Variational Inference (SVI), we identify orthotropic hyperelastic parameters from a single biaxial test; the overall workflow is summarized in Figure 1 and detailed in Section 2.2.
@@
Figure 1: Schematic overview of our unsupervised full-field stochastic variational inference framework for orthotropic hyperelastic tissue behavior. Quantification of the tissue’s microstructural organization, such as spatially varying fiber orientations, provides essential architectural information across the sample domain. In parallel, point-wise measurements of displacements $\hat{\bm{u}}$ and reaction forces $\hat{\bm{R}}^{\beta}$ are acquired under quasi-static biaxial loading using digital volume correlation. The sample geometry is discretized into a finite element mesh, enabling reconstruction of continuous deformation fields $\bm{F}$ from the measured data. A nonlinear orthotropic constitutive model $\psi$, parameterized by a set of cross correlated material parameters $\bm{\theta}$, serves as the mechanistic basis for inference. Given the reconstructed deformation and microstructural organization fields, the model predicts stress responses at the element level, which are used to compute internal and external nodal forces. Residuals are formulated through the weak form of the conservation of linear momentum, minimized pointwise for unconstrained degrees of freedom and in aggregate at boundaries with known reaction forces. The inverse problem is cast as a stochastic variational inference task, where a variational guide distribution $q(\bm{\theta})$ approximates the true posterior $\pi\left(\bm{\theta}\,|\,\hat{\bm{u}},\hat{\bm{R}}^{\beta}\right)$ by minimizing the Kullback-Leibler divergence $\mathrm{KL}\left(\pi\left(\bm{\theta}\,|\,\hat{\bm{u}},\hat{\bm{R}}^{\beta}\right)\,||\,q\left(\bm{\theta}\right)\right)$. Our physics-informed constitutive inference framework enables efficient and scalable inference from complex high dimensional full-field deformation datasets, accommodating nonlinear material behaviors and cross-correlated parameter effects.
@@ 2.1 EUCLID preliminaries
Consider a myocardial tissue specimen undergoing quasi-static deformation in a three-dimensional reference domain $\Omega\subset\mathbb{R}^{3}$. Dependent on the tissue slab’s extraction orientation protocol (Fig. 3), the underlying myocardial microstruture variation induces diverse and heterogeneous strain states, while the material itself is assumed to be homogeneous. Optionally, additional tissue manipulations, e.g. cutting a hole in the tissue slab or local cardiac ablation, can further promote a highly heterogeneous deformation state during biaxial tensile testing.
Boundary conditions are applied such that Dirichlet conditions are enforced on $\partial\Omega_{u}\subset\partial\Omega$ and Neumann conditions on the remainder, $\partial\Omega_{t}=\partial\Omega\setminus\partial\Omega_{u}$. For simplicity, our analysis focuses on displacement-controlled loading (i.e., Dirichlet boundary conditions), while noting that applied forces in load-controlled scenarios are equivalent to reaction forces under displacement control. The dataset comprises of $n_{t}$ snapshots of displacement measurements,
$\mathcal{U}=\left\{\bm{u}^{a,t}\in\mathbb{R}^{3}:a=1,\dots,n_{n};\ t=1,\dots,n_{t}\right\},$ (1)
recorded at $n_{n}$ reference points
$\mathcal{X}=\left\{\bm{X}^{a}\in\Omega:a=1,\dots,n_{n}\right\}.$ (2)
@@
$\left\{R^{\beta,t}:\beta=1,\dots,n_{\beta};\ t=1,\dots,n_{t}\right\}$ (3)
are measured at selected Dirichlet boundaries (e.g., using load cells). Since $n_{\beta}\ll n_{n}$, the force data is sparse compared to the dense displacement field. For brevity, the superscript $(\cdot)^{t}$ is omitted in the subsequent discussion, although the numerical procedure is applied independently to every snapshot. Given these limited measurements, the primary goal is to infer the a priori assumed constitutive model $\psi(\bm{F},\bm{\theta})$’s parameter set $\bm{\theta}$ governing the stress–strain response.
To reconstruct the displacement field, the reference domain $\mathcal{X}$ is discretized using linear tetrahedral elements (each with a single quadrature point at its barycenter),
@@
Here, $N^{a}:\Omega\to\mathbb{R}$ represents the shape function associated with node $\bm{X}^{a}$. We subsequently approximate the corresponding deformation gradient field as
@@
Following Flaschel et al. (2021) and Thakolkaran et al. (2022), we leverage the conservation of linear momentum to guide the learning of the constitutive model parameter set, which eliminates the need of stress labels. Assuming quasi-static loading conditions and negligible body forces, the weak form of the linear momentum balance in our reference domain $\Omega$ is given by
$\int_{\Omega}\bm{P}(\bm{F}):\nabla\bm{v}\,\mathrm{d}V-\int_{\partial\Omega_{t}}\hat{\bm{t}}\cdot\bm{v}\,\mathrm{d}S=0\quad\forall\ \text{admissible }\bm{v}\,$, (6)
where $\hat{\bm{t}}$ denotes the prescribed traction acting on $\partial\Omega_{t}$ and $\bm{v}$ is a test function that is sufficiently regular and vanishes on the Dirichlet boundary $\partial\Omega_{u}$. We prefer the weak formulation over the strong form, as it avoids the need for second order derivatives with respect to the displacement-field, which are sensitive to measurement noise.
Assuming no dissipative energy losses within the material and rewriting the Clausius–Duhem entropy inequality Planck (1897) following the Coleman and Noll principle Coleman and Noll (1959); Peirlinck et al. (2024), we get
@@
as the constitutive relation between the first Piola-Kirchoff stress $\bm{P}$ and the deformation gradient $\bm{F}$ defined through the free energy function $\psi$ and parametrized using the constitutive parameter set $\bm{\theta}$.
Let $\mathcal{D}=\{(a,i):a=1,\dots,n_{n};\ i=1,2,3\}$ denote all the displacement degrees of freedom in our domain $\Omega$, which we partition into $\mathcal{D}^{\text{free}}$, the unconstrained degrees of freedom, and $\mathcal{D}^{\text{fix}}_{\beta}$ (with $\beta=1,\dots,n_{\beta}$) denoting the degrees of freedom under Dirichlet constraints that contribute to the observed reaction force $R^{\beta}$.
Approximating the test function $\bm{v}$ by
@@
$\sum_{a=1}^{n_{n}}v_{i}^{a}f_{i}^{a}=0\,,\quad\text{with}\quad f_{i}^{a}=\underbrace{\int_{\Omega}P_{ij}\,\nabla_{j}N^{a}\,\mathrm{d}V}_{\text{internal force}}-\underbrace{\int_{\partial\Omega_{t}}\hat{t}_{i}N^{a}\,\mathrm{d}S}_{\text{external force}}\,.$ (9)
Here, the integrals are computed using numerical quadrature over the same discretization and mesh as (4) resulting in a Bubnov-Galerkin form. Since the test functions are arbitrary, the force residual must vanish at every unconstrained degree of freedom,
@@
At the fixed degrees of freedom, the internal and external forces are counteracted by the reaction force imposed by the Dirichlet constraints. Since point-wise reaction forces cannot be measured experimentally, they are assumed to be unavailable. Instead, only the total reaction forces integrated over the boundary segments are known. As a result, the global force balance for each measured reaction force is expressed as
@@
where the summation is carried out over all point-wise forces associated with the degrees of freedom in the $\beta^{\text{th}}$ Dirichlet constraint, denoted as $\mathcal{D}^{\text{fix}}_{\beta}$. As noted earlier, the superscript $(\cdot)^{t}$ has been omitted for brevity, but the above force balance conditions apply to all data snapshots at $t=1,\dots,n_{t}$.
Together, (10) and (11) constitute a system of equations, which defines the inverse problem of constitutive model inference (Flaschel et al., 2021; Thakolkaran et al., 2022, 2025) as the minimization of the force balance residuals with respect to the trainable parameter set $\bm{\theta}$ of the free energy function $\psi_{\bm{}}$ (7):
$\bm{\hat{\theta}}\leftarrow\arg\min_{\bm{\theta}}\sum_{t=1}^{n_{t}}\Bigg[\underbrace{\sum_{(a,i)\ \in\ \mathcal{D}^{\text{free}}}\left(f_{i}^{a,t}\right)^{2}}_{\text{free degrees of freedom}}+\underbrace{\sum_{\beta=1}^{n_{\beta}}\Bigl(R^{\beta,t}-\sum_{(a,i)\ \in\ \mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a,t}\Bigr)^{2}}_{\text{fixed degrees of freedom}}\Bigg]\,$, (12)

### v2

@@ Abstract
Cardiac muscle tissue exhibits highly non-linear hyperelastic and orthotropic material behavior during passive deformation. Traditional constitutive identification protocols therefore combine multiple loading modes and typically require multiple specimens and substantial handling. In soft living tissues, such protocols are challenged by inter- and intra-sample variability and by manipulation-induced alterations of mechanical response, which can bias inverse calibration. In this work we exploit spatially heterogeneous full-field kinematics as an information-rich alternative to multimodal testing. We adapt EUCLID, an unsupervised method for the automated discovery of constitutive models, towards Bayesian parameter inference for highly nonlinear, orthotropic constitutive models. Using synthetic myocardial tissue slabs, we demonstrate that a single heterogeneous biaxial experiment, combined with sparse reaction-force measurements, enables robust recovery of Holzapfel–Ogden parameters with quantified uncertainty, across multiple noise levels. The inferred responses agree closely with ground-truth simulations and yield credible intervals that reflect the impact of measurement noise on orthotropic material model inference. Our work supports single-shot, uncertainty-aware characterization of nonlinear orthotropic material models from a single biaxial test, reducing sample demand and experimental manipulation.
@@
material model inference , full-field kinematics , uncertainty quantification , anisotropic hyperelasticity , experimental tissue testing , Bayesian inference
@@ 1 Introduction
Soft biological tissues exhibit complex mechanical responses due to their fibrous, anisotropic microstructure. Accurately characterizing this behavior is essential for constructing predictive constitutive models that support simulation-driven design, diagnostics, and treatment planning [1, 2]. However, unlike engineered materials, soft tissues are subject to considerable biological variability, both within a patient, across a patient cohort, and across experimental studies [3, 4, 5, 6, 7]. This intra- and inter-sample variability is even more pronounced in the myocardium, where strong spatial gradients in myofiber orientation and laminar architecture govern regional mechanical function [8, 9]. Here, we investigate the impact that the underlying microstructural organization variability has on the mechanical behavior of a single sample.
@@
Traditional approaches to constitutive modeling of such tissues typically rely on multiple mechanical tests across different deformation modes including uniaxial tension or compression, biaxial tension, or shear testing [10, 11, 12]. These protocols often require combining data from different specimens and, as such, ignore and smooth out intrinsic intra- and inter-sample variability into an average response [13]. Moreover, these protocols involve strong assumptions, such as spatially homogeneous stress and strain fields within the region of interest [14, 15, 16]. Such assumptions are problematic for living tissues, where spatial heterogeneity is the norm, and the mechanical response may depend on subtle differences in fiber orientation, geometry, or mounting. In addition, invasive sample manipulations, such as cutting, clamping, and gluing, can alter the intrinsic mechanical behavior that these tests seek to measure [17, 18, 19].
@@
These challenges have prompted the development of single-sample, single-experiment methodologies that aim to infer full constitutive behavior from a single loading protocol. This shift has been enabled by advances in full-field deformation measurements, such as digital image correlation and imaging-based 3D kinematics [20, 21, 22]. When applied to samples undergoing spatially heterogeneous deformation, such data enable the inference of multiple material parameters simultaneously, reducing the need for multiple loading modes. Indeed, unsupervised identification techniques, such as the virtual fields method (VFM) [23, 24], finite element model updating (FEMU) [25, 26, 27], variational systems identification [20], the equilibrium gap method (EGM) [28, 29], and EUCLID [30, 31, 32, 33], demonstrate that heterogeneous strain states, once seen as a nuisance, can be leveraged as rich sources of information, obviating the need for homogenizing assumptions. In this study, we assess the extent to which a simple biaxial tensile testing protocol on a single myocardial tissue slab can support identification of the underlying orthotropic behavior.
@@
Yet, a key limitation persists in most current workflows: the lack of uncertainty quantification. Material parameters are frequently reported as point estimates, with little accounting for the inherent measurement noise or biological variability. In the context of myocardial modeling, where even adjacent samples can differ markedly in stiffness and fiber orientation [10, 34], such deterministic inference may be misleading [35]. Recent work has begun to address this gap through Bayesian approaches to inverse modeling, which offer not only best-fit estimates but also posterior distributions reflecting uncertainty in data and model form [36, 32, 37, 38, 39, 40, 41]. Within the unsupervised setting, Bayesian-EUCLID forms an interesting contribution towards the inference of myocardial tissue behavior from full-field displacement fields and force measurements. In prior work, Joshi et al. demonstrated the feasibility of using Gibbs sampling for full-field inference in two-dimensional problems [32], where the material response could be inferred from a linear system of equations for the unknown material parameters in a rich constitutive feature library.
In contrast, the setting of this study involves two distinct and compounding challenges. First, the underlying constitutive models are highly nonlinear [42, 14, 16], with strongly cross-correlated parameters [15, 43] that give rise to complex posterior landscapes. Second, we transition to three-dimensional full-field data. While the latter provides richer information, it also substantially increases the computational cost of evaluating the likelihood function. These factors independently and jointly limit the practical utility of Markov Chain Monte Carlo (MCMC) methods, including Gibbs sampling, due to the high cost of repeated likelihood evaluations and the poor mixing behavior in high-dimensional, correlated spaces. These limitations motivates adapting the Bayesian EUCLID framework with more scalable inference strategies, such as Stochastic Variational Inference (SVI), which approximates posterior distributions through optimization rather than sampling, enabling tractable inference in settings with both complex models and large data volumes.
@@
Towards these goals, this study presents an unsupervised, full-field, Bayesian framework for the inference of orthotropic hyperelastic material parameters from a single heterogeneous biaxial tensile test. Our approach builds on the Bayesian-EUCLID framework [32], extended via SVI to capture parametric uncertainty in a highly nonlinear, a priori chosen constitutive model form. The myocardium is modeled using the well-established Holzapfel–Ogden constitutive law [14, 44], and heterogeneity in the deformation field is introduced either through microstructural variation (i.e., intrinsic fiber-sheet-interlaminar architectural heterogeneity), geometric tissue sample modification, or both. We demonstrate that the spatial richness of the heterogeneous displacement field combined with reaction force data enables robust parameter inference with uncertainty quantification from a single test. We additionally validate the inferred material models using unseen invariant values and assess sensitivity to displacement noise. Overall, this work highlights the potential of full-field, unsupervised, and uncertainty-aware approaches to redefine how we efficiently characterize complex biological tissues, taking into account minimal tissue availability and minimizing tissue manipulation.
@@ 2 Unsupervised Stochastic Variational constitutive parameter inference
Conventional approaches for characterizing myocardial material behavior typically rely on spatially averaged measurements and simplifying assumptions of spatially homogeneous microstructural organization. However, this intrinsic homogenization obscures the spatial heterogeneity present within a tissue sample. To address this, we infer the constitutive parameter vector in an unsupervised setting from experimentally accessible data: (i) global reaction forces, (ii) full-field deformation, and (iii) microstructural organization across the slab. In this work, we assume digital volume correlation (DVC) data derived from concomitant through-thickness imaging, such as ultrasound [45], Magnetic Resonance Imaging [46], or X-Ray Imaging [47] to obtain three-dimensional deformation fields. In parallel, microstructural maps can be obtained via SHG or micro-CT [48, 12, 49]. Among established inverse constitutive identification strategies such as FEMU, VFM, and EGM, we build on EUCLID [30]. EUCLID formulates the inverse problem via the weak form of momentum balance (see Section 2.1). Here, we adapt it to a fully three-dimensional setting with nonlinear orthotropic materials and recast the classic deterministic parameter optimization into an unsupervised Bayesian inference problem. Leveraging Stochastic Variational Inference (SVI), we identify orthotropic hyperelastic parameters from a single biaxial test; the overall workflow is summarized in Figure 1 and detailed in Section 2.2.
@@
Figure 1: Schematic overview of our unsupervised full-field stochastic variational inference framework for orthotropic hyperelastic tissue behavior. Quantification of the tissue’s microstructural organization, such as spatially varying fiber orientations, provides essential architectural information across the sample domain. In parallel, point-wise measurements of displacements $\bm{u}$ and reaction forces $\bm{R}^{\beta}$ are acquired under quasi-static biaxial loading using digital volume correlation. The sample geometry is discretized into a finite element mesh, enabling reconstruction of continuous deformation fields $\bm{F}$ from the measured data. A nonlinear orthotropic constitutive model $\psi$, parameterized by a set of cross correlated material parameters $\bm{\theta}$, serves as the mechanistic basis for inference. Given the reconstructed deformation and microstructural organization fields, the model predicts stress responses at the element level, which are used to compute internal and external nodal forces. Residuals are formulated through the weak form of the conservation of linear momentum, minimized pointwise for unconstrained degrees of freedom and in aggregate at boundaries with known reaction forces. The inverse problem is cast as a stochastic variational inference task, where a variational guide distribution $q(\bm{\theta})$ approximates the true posterior $\pi\left(\bm{\theta}\,|\,\bm{u},\bm{R}^{\beta}\right)$ by minimizing the Kullback-Leibler divergence $\mathrm{KL}\left(q\left(\bm{\theta}\right)\,||\,\pi\left(\bm{\theta}\,|\,\bm{u},\bm{R}^{\beta}\right)\right)$. Our physics-informed constitutive inference framework enables efficient and scalable inference from complex high dimensional full-field deformation datasets, accommodating nonlinear material behaviors and cross-correlated parameter effects.
@@ 2.1 EUCLID preliminaries
We first define the kinematic and force data available from a single experiment, and then derive the equilibrium residuals that underpin deterministic parameter identification. Consider a myocardial tissue specimen undergoing quasi-static deformation in a three-dimensional reference domain $\Omega\subset\mathbb{R}^{3}$. Depending on the tissue slab’s extraction orientation protocol (see Section  3), the underlying myocardial microstructure variation induces diverse and heterogeneous strain states. In this work, this heterogeneity enters through spatially varying structural fields, while the constitutive parameter vector $\bm{\theta}$ is assumed to be spatially homogeneous and thus constant over $\Omega$. This separation attributes spatial variability to the structural inputs while keeping the inverse problem focused on a single parameter vector. Optionally, additional tissue manipulations, e.g., cutting a hole in the tissue slab or local cardiac ablation, can further promote a highly heterogeneous deformation state during biaxial tensile testing.
Boundary conditions are applied such that Dirichlet conditions are enforced on $\partial\Omega_{u}\subset\partial\Omega$ and Neumann conditions on the remainder, $\partial\Omega_{t}=\partial\Omega\setminus\partial\Omega_{u}$. For simplicity, our analysis focuses on displacement-controlled loading (i.e., Dirichlet boundary conditions), while noting that applied forces in load-controlled scenarios are equivalent to reaction forces under displacement control. The dataset comprises $n_{t}$ snapshots of displacement measurements,
$\left\{\bm{u}^{a,t}\in\mathbb{R}^{3}:a=1,\dots,n_{n};\ t=1,\dots,n_{t}\right\},$ (1)
recorded at $n_{n}$ reference points comprising the measurement node set:
$\left\{\bm{X}^{a}\in\Omega:a=1,\dots,n_{n}\right\}.$ (2)
@@
$\left\{R^{\beta,t}\in\mathbb{R}:\beta=1,\dots,n_{\beta};\ t=1,\dots,n_{t}\right\}$ (3)
are measured at selected Dirichlet boundaries (e.g., using load cells). Since $n_{\beta}\ll n_{n}$, the reaction force data are sparse compared to the dense displacement field. For brevity, the superscript $(\cdot)^{t}$ is omitted in the subsequent discussion, although the numerical procedure is applied independently to every snapshot. Given $\left\{\bm{u}^{a}\right\}$ and $\left\{R^{\beta}\right\}$, our goal is to infer the parameter set $\bm{\theta}$ of an a priori assumed constitutive model $\psi(\bm{F},\bm{\theta})$.
To connect the measurements to equilibrium, we first reconstruct a continuous kinematic field and its deformation gradient. The reference domain $\Omega$ is discretized using linear tetrahedral elements, evaluated with a single quadrature point at its barycenter, yielding
@@
Here, $N^{a}:\Omega\to\mathbb{R}$ represents the shape function associated with the measurement node $\bm{X}^{a}$. We subsequently approximate the corresponding deformation gradient field as
@@
Assuming hyperelasticity, i.e., no dissipative energy losses within the material, the first Piola Kirchhoff stress follows as
@@
as the constitutive relation between the first Piola Kirchhoff stress $\bm{P}$ and the deformation gradient $\bm{F}$ defined through the free energy function $\psi$ and parameterized using the constitutive parameter set $\bm{\theta}$ [50, 51, 52]. In particular, $\bm{P}$ depends on the unknown parameters through $\psi(\bm{F},\bm{\theta})$.
We leverage the conservation of linear momentum to guide the learning of the constitutive model parameter set, which eliminates the need for stress labels [30, 31]. Assuming quasi-static loading conditions and negligible body forces, the weak form of the linear momentum balance in our reference domain $\Omega$ is given by
$\int_{\Omega}\bm{P}(\bm{F},\bm{\theta}):\nabla\bm{v}\,\mathrm{d}V-\int_{\partial\Omega_{t}}\bar{\bm{t}}\cdot\bm{v}\,\mathrm{d}S=0\quad\forall\ \text{admissible }\bm{v}\,$, (7)
where $\bar{\bm{t}}$ denotes the prescribed traction acting on $\partial\Omega_{t}$ per unit reference area and $\bm{v}$ is a test function that is sufficiently regular and vanishes on the Dirichlet boundary $\partial\Omega_{u}$. We prefer the weak formulation over the strong form, as it avoids the need for second order derivatives with respect to the displacement field, which are sensitive to measurement noise. In our displacement-controlled setting, non-instrumented free surfaces are typically traction free (so $\bar{\bm{t}}=0$ there). Measured boundary forces enter through reactions on $\partial\Omega_{u}$, which we incorporate below.
Let $\mathcal{D}=\{(a,i):a=1,\dots,n_{n};\ i=1,2,3\}$ denote all the displacement degrees of freedom in our domain $\Omega$, which we partition into $\mathcal{D}^{\text{free}}$, the unconstrained degrees of freedom, and $\mathcal{D}^{\text{fix}}_{\beta}$ (with $\beta=1,\dots,n_{\beta}$) denoting the degrees of freedom under Dirichlet constraints that contribute to the observed reaction force $R^{\beta}$. Using a Bubnov–Galerkin discretization, we approximate the admissible test function $\bm{v}$ as
@@
$\sum_{a=1}^{n_{n}}v_{i}^{a}f_{i}^{a}=0\,,\quad\text{with}\quad f_{i}^{a}=\underbrace{\int_{\Omega}P_{ij}\,\nabla_{j}N^{a}\,\mathrm{d}V}_{\text{internal force}}-\underbrace{\int_{\partial\Omega_{t}}\bar{t}_{i}N^{a}\,\mathrm{d}S}_{\text{external force}}\,.$ (9)
The integrals are evaluated by numerical quadrature on the same discretization as (4). Since the test functions are arbitrary on $\mathcal{D}^{\text{free}}$, the force residual must vanish at every unconstrained degree of freedom,
@@
At the fixed degrees of freedom, the internal and external forces are counteracted by the reaction force imposed by the Dirichlet constraints. Because point-wise reactions are not accessible experimentally, we treat them as unavailable. Instead, we use the measured reaction forces integrated over instrumented boundary segments, which yields the global balance constraint
@@
where the summation is carried out over all point-wise forces associated with the degrees of freedom in the $\beta^{\text{th}}$ Dirichlet constraint, denoted as $\mathcal{D}^{\text{fix}}_{\beta}$.
Together, (10) and (11) constitute a system of equations that defines the inverse problem of constitutive model inference [30, 31, 33]. We deterministically solve for $\bm{\theta}^{\star}$ by minimizing the equilibrium residuals over all snapshots:
$\bm{\theta}^{\star}\leftarrow\arg\min_{\bm{\theta}}\sum_{t=1}^{n_{t}}\Bigg[\underbrace{\sum_{(a,i)\ \in\ \mathcal{D}^{\text{free}}}\left(f_{i}^{a,t}\right)^{2}}_{\text{free degrees of freedom}}+\underbrace{\sum_{\beta=1}^{n_{\beta}}\Bigl(R^{\beta,t}-\sum_{(a,i)\ \in\ \mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a,t}\Bigr)^{2}}_{\text{fixed degrees of freedom}}\Bigg]\,$, (12)